OpenStudy (anonymous):

What are the sine, cosine, and tangent of radians (7pi)/(4)

OpenStudy (anonymous):

\[\frac{ 7\pi }{ 4 }\]

OpenStudy (anonymous):

@ganeshie8 @Hero @jdoe0001 @dumbcow

OpenStudy (dumbcow):

use a unit circle cos is like the "x" coordinate sin is like the "y" coordinate tan = sin/cos